OpenStudy (anonymous):

differential equations help!! has anyone taken the class recently???? I really need help on these problems

OpenStudy (anonymous):

\[xy'''-3y'+e^xy=x^2-1\] \[y(-2)=1, y'(-2)=0, y''(-2)=2\] the answer is (-infinity,0)

OpenStudy (anonymous):

In Problems 1–6, determine the largest interval (a,b) for which Theorem 1 guarantees the existence of a unique solution on (a,b) to the given initial value problem

OpenStudy (anonymous):

@Mertsj @phi @robtobey if you guys could help or mention someone that could it would be appreciated. im looking at the book now to see how it's done......

OpenStudy (phi):

what is Thm 1 ?

OpenStudy (anonymous):

http://prntscr.com/1hqnvh

OpenStudy (anonymous):

@phi there is theorem 1

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (anonymous):

@oldrin.bataku don't worry about it friend already got this one...just took about 2 hours for someone to finally help lol

OpenStudy (anonymous):

I want to know, ignore him,please tell me how to solve it.

OpenStudy (anonymous):

$$xy'''-3y'+e^xy=x^2-1$$Observe that theorem 1 considers IVPs in standard form, so divide thru by \(x\):$$y'''-\frac3xy'+\frac{e^x}xy=\frac{x^2-1}x$$We're interested in intervals of continuity, and clearly our coefficient functions are all continuous on \((-\infty,0),(0,\infty)\) (because of the poles at \(x=0\)). We're interested in the largest interval of continuity containing the initial point used for our initial values, here \(x_0=-2\), hence we pick \((-\infty,0)\).

OpenStudy (anonymous):

sorry @rperez36 I was eating. @looser66 see above

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

@looser66 that's it :-p theorem 1 tells us that the IVP has a unique solution on any interval on which our coefficient functions are continuous and we picked the largest one that contains \(-2\)